in triangle ABC angle b is 20 degrees more than angle a. The sun of angle a and b is 20 degrees more than angle c. set up and solve a system of equations to find angles a,b, and c
Answer by ReadingBoosters(3246)

a + b + c = 180
b = 20 + a, which is the same as b - 20 = a
a + b = 20 + c
Substitute the second equation in the third equation
a + 20 + a = 20 + c
2a + 20 = 20 + c
2a + 20 - 20 = 20 + c - 20
2a = c
Substitute into first equation
a + 20 + a + 2a = 180
4a + 20 = 180
4a + 20 - 20 = 180 - 20
4a = 160
4a/4 = 160/4
a = 40
Solve for b, where b = 20+a
b = 60
Solve for c, where c = 2a
c = 80